OpenStudy (anonymous):

(1/R_1)+(1/R_2)=(1/R_eq) solve for R_2

OpenStudy (anonymous):

written like this\[(1/R_1)+(1/R_2)=(1/R_eq) solve for R_2\]

OpenStudy (phi):

one good first step is subtract 1/R1 from both sides you get 1/R2 = 1/Req - 1/R1 the next step is find a common denominator for the stuff on the right side and do the subtraction can you do that ?